# The Russian-focused embedders’ exploration: ruMTEB benchmark and Russian embedding model design

### v1

@@ Abstract
Embedding models play a crucial role in Natural Language Processing (NLP) by creating text embeddings used in various tasks such as information retrieval and assessing semantic text similarity. This paper focuses on research related to embedding models in the Russian language. It introduces a new Russian-focused embedding model called ru-en-RoSBERTa and the ruMTEB benchmark, the Russian version extending the Massive Text Embedding Benchmark (MTEB). Our benchmark includes seven categories of tasks, such as semantic textual similarity, text classification, reranking, and retrieval. The research also assesses a representative set of Russian and multilingual models on the proposed benchmark. The findings indicate that the new model achieves results that are on par with state-of-the-art models in Russian. We release the model ru-en-RoSBERTa, and the ruMTEB framework comes with open-source code, integration into the original framework and a public leaderboard 111https://huggingface.co/spaces/mteb/leaderboard.
@@ 1 Introduction
Text embeddings play an important role in many Natural Language Processing (NLP) tasks, from clustering to semantic textual similarity (STS) and information retrieval (IR). The community has addressed this demand by releasing several powerful text embedding models (or embedders) Wang et al. (2024, 2023a); Chen et al. (2024). However, there is still a lack of such embedders developed specifically for the Russian language. The most popular Russian-oriented models, such as rubert-tiny2 222https://huggingface.co/cointegrated/rubert-tiny2, SBERT333https://huggingface.co/ai-forever/sbert_large_nlu_ru, and SBERT444https://huggingface.co/ai-forever/sbert_large_mt_nlu_ru, have been released several years ago and thus do not include modern data in their training corpora. The latest models are based on an outdated version of the ruBERT Zmitrovich et al. (2023) model as a backbone. Moreover, being monolingual, they can not profit from knowledge transfer between languages.
@@
we present the Russian version of MTEB and release 17 new Russian datasets for text embedding evaluation 777https://huggingface.co/collections/ai-forever/ru-mteb-6650a6a6708dc5107a9e0ba3, which form the benchmark backbone;
@@ 2.2 Text Embedding Benchmarks
Model evaluation has always played an inevitable role in NLP progress. Starting from GLUE Wang et al. (2018) and SuperGLUE Wang et al. (2019) benchmarks have been a standard model evaluation method. As for text embedding representation evaluation, it has been in focus for many years and was commonly evaluated on a STS, for which yearly released SemEval Agirre et al. (2016); Cer et al. (2017); Chen et al. (2022)888https://semeval.github.io/ datasets were commonly used. Being a single dataset inevitably limits the SemEval expressivity. Following the same approach, SentEval Conneau and Kiela (2018), which focuses on classifier models on top of embedding, overcomes this limitation by aggregating multiple STS datasets. Still, it lacks the evaluation instruments for the suitability of embedding for retrieval or clustering tasks. Due to the inefficiency of a single STS evaluation, USEB Wang et al. (2021), focusing on reranking tasks, and BEIR Thakur et al. (2021), aimed at zero-shot information retrieval evaluation, were created. SciRepEval Singh et al. (2023), a multi-format benchmark for scientific document representations, includes 24 realistic tasks across four formats: classification, regression, ranking, and search. Finally, uniting and unifying all main classes of the embedding tasks MTEB Muennighoff et al. (2023) has been proposed and is now considered a multilingual text embedding evaluation standard. Moreover, its approach was also adopted and recreated for the Scandinavian languages in the SEB Enevoldsen et al. (2024) and Chinese in the C-MTEB Xiao et al. (2023a)999https://huggingface.co/C-MTEB benchmarks.
@@ 3.1 Benchmark Structure and Evaluation Methodology
Reranking. Inputs are a query and a list of reference texts (relevant and irrelevant). The goal is to correctly rank these texts according to their relevance to the query. For evaluation, the texts for each query are ranked by the cosine similarity between the query embedding and the embedding of the given texts. The obtained ranking is scored with MAP@k () for each query and averaged over all queries.
@@ 3.2 Benchmark Tasks
We took the datasets based on the original MTEB without any changes. For the datasets adapted from other Russian sets, we performed data cleaning and automatic filtering, where necessary, removed duplicates, and formatted them in the MTEB format. The main dataset information and their statics are given in Table 1, and the detailed task descriptions and preprocessing for the new sets are in Appendix A.3.
@@ 4.1 Training Data
Following previous work Wang et al. (2022a); Li et al. (2023); Nussbaum et al. (2024), we use publicly available data, high-quality and synthetic datasets to create training pairs (see Appendix A.1.1 for the full training list) 101010To avoid potential data leakage we use only the training parts of all the sets., which, for experiment purpose (see Section 5), we divide into four groups described below.
@@
Additional Synthetic Datasets. The group includes Query2doc MS-MARCO Wang et al. (2023b), DINO-STS-x1x2 Schick and Schütze (2021), RuHNP Malashenko et al. (2024a), entailment and contradiction pairs from RuWANLI Malashenko et al. (2024b), and a sample of generated pairs by ruT5-base111111https://huggingface.co/ai-forever/ruT5-base model from WikiOmnia Pisarevskaya and Shavrina (2022). We do not change the data content and use the datasets as is.
@@ 4.2 Base Model and English Language Adaptation
Since we focus on the Russian language, we use ruRoBERTa 121212https://huggingface.co/ai-forever/ruRoberta-large Zmitrovich et al. (2023), which has the highest scores on the classic Russian SuperGLUE Shavrina et al. (2020) benchmark among the models of its size. In addition, we adapt it to the English language, allowing knowledge transfer from this high-resource language (see Section 5 for the corresponding experiments).
We extend the original ruRoBERTa tokenizer with tokens from RoBERTa 131313https://huggingface.co/FacebookAI/roberta-large Liu et al. (2019). To learn new token embeddings, we train the model using Masked Language Modeling (MLM) objective Devlin et al. (2018). We use the same hyperparameters as in the ruRoBERTa and the batch size of 1024. We use unique training texts from Section 4.1 and train for one epoch (~11k steps).
The whole process takes one day on two A100 80GB cards. To reduce the effect of catastrophic forgetting Kirkpatrick et al. (2017), we merge encoder layers using spherical linear interpolation (SLERP) algorithm141414https://gist.github.com/dvschultz/3af50c40df002da3b751efab1daddf2c with the factor of 0.25 to the original model.
In our work, we refer to the obtained model version with the extended vocabulary as ru-en-RoBERTa.
@@ 5 Training Procedure Analysis
This section describes experiments we conducted to determine the final training pipeline. We used basic Russian, English, and additional synthetic datasets, the training approach described in Section 4 unless otherwise specified, and the full ruMTEB version for evaluation. Further findings are given in Appendix A.6.
@@ 5.2 English Language Adaptation
| Increase hard-negatives group | 62.7 | 55.6 | 38.4 | 60.6 | 63.1 | 65.7 | 73.5 | 59.9 |
@@ 5.3 Training Examples
Hard-negatives. To study whether adding more hard-negatives Ren et al. (2021) is beneficial, we increased their number to 15 and reduced per device batch size to 64, maintaining the total number of negative examples. To keep the same number of steps, we apply gradient accumulation. Similarly to Nussbaum et al. (2024), we found that despite processing almost twice as many texts, the results did not improve.
@@ 5.4 Training Objective
| rubert-tiny2 | 52.1 | 38.7 | 29.4 | 51.9 | 31.0 | 8.9 | 61.9 | 39.1 |
| SBERT | 57.2 | 51.9 | 31.9 | 50.2 | 32.8 | 8.5 | 57.2 | 41.4 |
| SBERT | 57.5 | 52.6 | 32.6 | 52.0 | 40.6 | 19.1 | 64.4 | 45.6 |
| mE5 | 56.5 | 51.3 | 31.4 | 55.1 | 65.3 | 65.8 | 69.5 | 56.4 |
| mE5 | 58.2 | 50.3 | 32.9 | 55.0 | 66.2 | 67.1 | 70.2 | 57.1 |
| mE5 | 60.9 | 52.5 | 35.3 | 58.4 | 69.6 | 74.0 | 71.6 | 60.4 |
| BGE-M3 | 60.3 | 52.2 | 34.5 | 60.6 | 69.7 | 74.8 | 73.7 | 60.8 |
| ru-en-RoSBERTa | 62.7 | 56.1 | 38.9 | 60.8 | 63.9 | 66.5 | 74.0 | 60.4 |
| mE5 | 66.1 | 63.3 | 40.4 | 63.5 | 69.8 | 75.1 | 74.9 | 64.7 |
| E5 | 68.8 | 64.0 | 42.2 | 60.5 | 69.9 | 74.1 | 74.4 | 64.9 |
@@ 6 Evaluation
We evaluate all models in the same environments and scenarios by the procedure described in 3.1. We use MTEB framework151515https://github.com/embeddings-benchmark/mteb/tree/1.12.75 for evaluation where we integrated evaluation on the new ruMTEB tasks 161616https://github.com/embeddings-benchmark/mteb/pull/815171717https://github.com/embeddings-benchmark/mteb/pull/881.
@@ 7 Results
Table 5 shows model scores averaged within the task category, and detailed results of the task-wise model evaluation are in Appendix A.5.
@@ 9 Limitation
Lack of evaluation in English. As this study focuses on Russian, we do not perform a comprehensive evaluation of ru-en-RoSBERTa in English.
@@
Datasets. The collaborative nature and aggregation of the existing sets in the benchmark make it challenging to ensure uniformly high data quality across all tasks.
For all benchmark datasets we checked the licenses and filtered the datasets. Unfortunately, despite the joint effort, some tasks still possess errors (e.g., incorrect labels for some examples, grammatical errors, surplus technical symbols, etc.). Moreover, there may still be biases in the data across different domains and sources, and there is still a need to extend tasks in some categories. We encourage researchers to collaborate further to fill the gaps and ensure a more comprehensive and balanced language and task representation in the benchmark.
@@ 10 Ethical Considerations
AI-assistants Help. We improve and proofread the text of this paper using Grammarly181818https://app.grammarly.com/ to correct grammatical, spelling, and style errors and paraphrasing sentences. Thus, some segments of our publication can be potentially detected as AI-generated, AI-edited, or human-AI-generated.
@@ A.1.1 Training Data Information
| Basic Russian Datasets 191919For non-cited datasets please refer to A.1.2 | |||
@@ A.1.2 Data Filtration Details
We apply the following steps to the basic Russian datasets. First, texts longer than 500 tokens (ruRoBERTa-large202020https://huggingface.co/ai-forever/ruRoberta-large tokenizer is used) are filtered out. A small number of tokens is reserved for instructions or prefixes. Pairs from YandexQ212121https://huggingface.co/datasets/IlyaGusev/yandex_q_full, Pikabu222222https://huggingface.co/datasets/IlyaGusev/pikabu, StackOverflow232323https://huggingface.co/datasets/IlyaGusev/ru_stackoverflow, Habr242424https://huggingface.co/datasets/IlyaGusev/habr and Habr QnA252525https://huggingface.co/datasets/its5Q/habr_qna are filtered by content popularity (e.g. views, ratings, votes). Cosine similarity obtained from LaBSE Feng et al. (2022) is applied to filter NewsCommentary and MultiParaCrawl. We filter pairs from paraphrase-NMT-Leipzig262626https://huggingface.co/datasets/cointegrated/ru-paraphrase-NMT-Leipzig by p_good score (equivalent meaning). The XNLI is formed from entailment (relevant document) and contradiction (irrelevant negative) examples. For MIRACL, we use the title as the query and the first paragraphs (until we reach the token limit) as the document. We form pairs for Paraphrases 272727https://huggingface.co/datasets/inkoziev/paraphrases from paraphrases field, taking one as a query and the others as positive documents. The content of RuNews282828https://huggingface.co/datasets/IlyaGusev/ru_news is not changed. After exact match deduplication, the final training pairs for all datasets are randomly sampled from the remaining pairs.
@@ A.2 Model Training Details
We train the model in bf16 dtype with gradient checkpointing and use AdamW Loshchilov and Hutter (2017) with a learning rate of 1e-5 and weight decay of 0.01 for exactly one epoch, which is approximately 3700 steps, of which linear warmup is 200 steps. After fine-tuning, the SLERP merging is applied to the base model with a factor of 0.1.
We apply stratified sampling per device batch (mini-batch). Therefore, the global batch includes mini-batches consisting of pairs of different datasets. On the one hand, it becomes impossible to exchange negatives between devices and thus scale the number of in-batch negatives. On the other hand, this increases the diversity of sources in the global batch. Therefore, we do not apply the DisCo Chen et al. (2023) trick to exchange negative examples across devices. The batch size is 128 per device, giving 1024 documents per query. Context length is set to 512 for queries and documents Merrick et al. (2024).
Training is conducted on a single H100 node. We utilize codebase from BGE292929https://github.com/FlagOpen/FlagEmbedding and adapt it for our experiments. PyTorch’s expandable_segments helps us to mitigate fragmentation issues due to variable sequence length.

### v2

@@ Abstract
Embedding models play a crucial role in Natural Language Processing (NLP) by creating text embeddings used in various tasks such as information retrieval and assessing semantic text similarity. This paper focuses on research related to embedding models in the Russian language. It introduces a new Russian-focused embedding model called ru-en-RoSBERTa and the ruMTEB benchmark, the Russian version extending the Massive Text Embedding Benchmark (MTEB). Our benchmark includes seven categories of tasks, such as semantic textual similarity, text classification, reranking, and retrieval. The research also assesses a representative set of Russian and multilingual models on the proposed benchmark. The findings indicate that the new model achieves results that are on par with state-of-the-art models in Russian. We release the model ru-en-RoSBERTa, and the ruMTEB framework comes with open-source code, integration into the original framework and a public leaderboard.
@@ 1 Introduction
Text embeddings play an important role in many Natural Language Processing (NLP) tasks, from clustering to semantic textual similarity (STS) and information retrieval (IR). The community has addressed this demand by releasing several powerful text embedding models (or embedders) Wang et al. (2024, 2023a); Chen et al. (2024). However, there is still a lack of such embedders developed specifically for the Russian language. The most popular Russian-oriented models, such as rubert-tiny2 111https://huggingface.co/cointegrated/rubert-tiny2, SBERT222https://huggingface.co/ai-forever/sbert_large_nlu_ru, and SBERT333https://huggingface.co/ai-forever/sbert_large_mt_nlu_ru, have been released several years ago and thus do not include modern data in their training corpora. The latest models are based on an outdated version of the ruBERT Zmitrovich et al. (2023) 444https://huggingface.co/ai-forever/ruBert-large model as a backbone. Moreover, being monolingual, they can not profit from knowledge transfer between languages.
@@
we present the Russian version of MTEB and release 17 new Russian datasets for text embedding evaluation 777https://huggingface.co/collections/ai-forever/ru-mteb-6650a6a6708dc5107a9e0ba3, which form the benchmark backbone, and a public leaderboard 888https://huggingface.co/spaces/mteb/leaderboard;
@@ 2.2 Text Embedding Benchmarks
Model evaluation has always played an inevitable role in NLP progress. Starting from GLUE Wang et al. (2018) and SuperGLUE Wang et al. (2019) benchmarks have been a standard model evaluation method. As for text embedding representation evaluation, it has been in focus for many years and was commonly evaluated on a STS, for which yearly released SemEval Agirre et al. (2016); Cer et al. (2017); Chen et al. (2022)999https://semeval.github.io/ datasets were commonly used. Being a single dataset inevitably limits the SemEval expressivity. Following the same approach, SentEval Conneau and Kiela (2018), which focuses on classifier models on top of embedding, overcomes this limitation by aggregating multiple STS datasets. Still, it lacks the evaluation instruments for the suitability of embedding for retrieval or clustering tasks. Due to the inefficiency of a single STS evaluation, USEB Wang et al. (2021), focusing on reranking tasks, and BEIR Thakur et al. (2021), aimed at zero-shot information retrieval evaluation, were created. SciRepEval Singh et al. (2023), a multi-format benchmark for scientific document representations, includes 24 realistic tasks across four formats: classification, regression, ranking, and search. Finally, uniting and unifying all main classes of the embedding tasks MTEB Muennighoff et al. (2023) has been proposed and is now considered a multilingual text embedding evaluation standard. Moreover, its approach was also adopted and recreated for the Scandinavian languages in the SEB Enevoldsen et al. (2024) and Chinese in the C-MTEB Xiao et al. (2023a)101010https://huggingface.co/C-MTEB benchmarks.
@@ 3.1 Benchmark Structure and Evaluation Methodology
Reranking. Inputs are a query and a list of reference texts (relevant and irrelevant). The goal is to correctly rank these texts according to their relevance to the query. For evaluation, the texts for each query are ranked by the cosine similarity between the query embedding and the embedding of the given texts. The obtained ranking is scored with MAP@k ()111111The exception is MIRACLReranking which is evaluated using nDCG@10 following the original MTEB methodology. for each query and averaged over all queries.
@@ 3.2 Benchmark Tasks
We took the datasets based on the original MTEB without any changes. For the Russian community datasets, we selected only the tests with high-quality labeling, relying on the original publications. We performed data cleaning and automatic filtering where necessary, removed duplicates, manually verified small subsets of examples, and formatted them in the MTEB format. The main dataset information and their statics are given in Table 1, and the detailed task descriptions and preprocessing for the new sets are in Appendix A.3.
@@ 4.1 Training Data
Following previous work Wang et al. (2022a); Li et al. (2023); Nussbaum et al. (2024), we use publicly available data, high-quality and synthetic datasets to create training pairs (see Appendix A.1.1 for the full training list) 121212To avoid potential data leakage we use only the training parts of all the sets., which, for experiment purpose (see Section 5), we divide into four groups described below.
@@
Additional Synthetic Datasets. The group includes Query2doc MS-MARCO Wang et al. (2023b), DINO-STS-x1x2 Schick and Schütze (2021), RuHNP Malashenko et al. (2024a), entailment and contradiction pairs from RuWANLI Malashenko et al. (2024b), and a sample of generated pairs by ruT5-base131313https://huggingface.co/ai-forever/ruT5-base model from WikiOmnia Pisarevskaya and Shavrina (2022). We do not change the data content and use the datasets as is.
@@ 4.2 Base Model and English Language Adaptation
Since we focus on the Russian language, we use ruRoBERTa 141414https://huggingface.co/ai-forever/ruRoberta-large Zmitrovich et al. (2023), which has the highest scores on the classic Russian SuperGLUE Shavrina et al. (2020) benchmark among the models of its size. In addition, we adapt it to the English language, allowing knowledge transfer from this high-resource language (see Section 5 for the corresponding experiments).
We extend the original ruRoBERTa tokenizer with tokens from RoBERTa 151515https://huggingface.co/FacebookAI/roberta-large Liu et al. (2019). To learn new token embeddings, we train the model using Masked Language Modeling (MLM) objective Devlin et al. (2018). We use the same hyperparameters as in the ruRoBERTa and the batch size of 1024. We use unique training texts from Section 4.1 and train for one epoch (~11k steps).
The whole process takes one day on two A100 80GB cards. To reduce the effect of catastrophic forgetting Kirkpatrick et al. (2017), we merge encoder layers using spherical linear interpolation (SLERP) algorithm161616https://gist.github.com/dvschultz/3af50c40df002da3b751efab1daddf2c with the factor of 0.25 to the original model. In our work, we refer to the obtained model version with the extended vocabulary as ru-en-RoBERTa.
@@ 5 Training Procedure Analysis
This section describes experiments we conducted to determine the final training pipeline. We used basic Russian, English, and additional synthetic datasets, the training approach described in Section 4 unless otherwise specified, and the full ruMTEB version for evaluation. Details on the model configurations in these experiments are given in A.2.2 and further findings are in Appendix A.6.
@@ 5.2 English Language Adaptation
| Increase hard negatives group | 62.7 | 55.6 | 38.4 | 60.6 | 63.1 | 65.7 | 73.5 | 59.9 |
@@ 5.3 Training Examples
Hard negatives. To study whether adding more hard negatives Ren et al. (2021) is beneficial, we increased their number to 15 and reduced per device batch size to 64, maintaining the total number of negative examples. To keep the same number of steps, we apply gradient accumulation. Similarly to Nussbaum et al. (2024), we found that despite processing almost twice as many texts, the results did not improve.
@@ 5.4 Training Objective
| rubert-tiny2 | 52.17 | 39.12 | 29.45 | 51.87 | 30.95 | 8.89 | 61.60 | 42.22 |
| SBERT | 57.24 | 50.44 | 31.87 | 50.17 | 32.81 | 8.51 | 57.21 | 45.35 |
| SBERT | 57.52 | 51.29 | 32.67 | 51.97 | 40.56 | 19.13 | 64.40 | 48.72 |
| mE5 | 56.44 | 51.35 | 31.99 | 55.14 | 65.28 | 65.85 | 69.48 | 57.29 |
| mE5 | 58.26 | 50.27 | 33.65 | 54.98 | 66.24 | 67.14 | 70.16 | 58.34 |
| mE5 | 61.01 | 52.23 | 36.00 | 58.42 | 69.65 | 74.04 | 71.62 | 61.41 |
| BGE-M3 | 60.46 | 52.38 | 34.86 | 60.60 | 69.71 | 74.79 | 73.68 | 61.58 |
| ru-en-RoSBERTa | 62.74 | 56.06 | 38.88 | 60.79 | 63.89 | 66.52 | 73.97 | 61.77 |
| mE5 | 66.31 | 63.21 | 41.15 | 63.89 | 69.17 | 74.41 | 74.85 | 66.03 |
| E5 | 69.11 | 64.24 | 42.93 | 60.81 | 69.96 | 74.19 | 73.71 | 67.18 |
@@ 6 Evaluation
We evaluate all models in the same environments and scenarios by the procedure described in 3.1. We use MTEB framework171717https://github.com/embeddings-benchmark/mteb/tree/1.14.12 for evaluation where we integrated evaluation on the new ruMTEB tasks 181818https://github.com/embeddings-benchmark/mteb/pull/815191919https://github.com/embeddings-benchmark/mteb/pull/881.
@@ 7 Results
Table 5 shows model scores averaged within the task category, and detailed results of the task-wise model evaluation are in Appendix A.5202020This results are valid for 09.10.2024. Please, refer to the leaderboard at https://huggingface.co/spaces/mteb/leaderboard for the latest results..
@@ 9 Limitation
Lack of evaluation in English. In this work, we focus on the Russian language, and therefore, we do not conduct ru-en-RoSBERTa evaluation in English as this is beyond the scope of this work and quite resource-consuming. Nevertheless, we acknowledge that evaluating the model on the machine translation task or on the English data (e.g., the full MTEB benchmark) is valuable. We leave this to future work.
@@
Datasets. The collaborative nature and aggregation of the existing sets in the benchmark make it challenging to ensure uniformly high data quality across all tasks. For all benchmark datasets we checked the licenses and filtered the datasets. Unfortunately, despite the joint effort, some tasks still possess errors (e.g., incorrect labels for some examples, grammatical errors, surplus technical symbols, etc.). Moreover, there may still be biases in the data across different domains and sources, and there is still a need to extend tasks in some categories. We encourage researchers to collaborate further to fill the gaps and ensure a more comprehensive and balanced language and task representation in the benchmark.
@@ 10 Ethical Considerations
AI-assistants Help. We improve and proofread the text of this paper using Grammarly212121https://app.grammarly.com/ to correct grammatical, spelling, and style errors and paraphrasing sentences. Thus, some segments of our publication can be potentially detected as AI-generated, AI-edited, or human-AI-generated.
Acknowledgments
The authors express their sincere gratitude to all those who contributed to the success of the ruMTEB project. In particular, the authors would like to thank David Dale for his collaboration and valuable insights. The authors also acknowledge Ilya Gusev for his significant contributions to the dataset creation. Special thanks are extended to Roman Solomatin for his essential role in organizing the leaderboard.
@@ A.1.1 Training Data Information
| Basic Russian Datasets 222222For non-cited datasets please refer to A.1.2 | |||
@@ A.1.2 Data Filtration Details
We apply the following steps to the basic Russian datasets. First, texts longer than 500 tokens (ruRoBERTa-large232323https://huggingface.co/ai-forever/ruRoberta-large tokenizer is used) are filtered out. A small number of tokens is reserved for instructions or prefixes. Pairs from YandexQ242424https://huggingface.co/datasets/IlyaGusev/yandex_q_full, Pikabu252525https://huggingface.co/datasets/IlyaGusev/pikabu, StackOverflow262626https://huggingface.co/datasets/IlyaGusev/ru_stackoverflow, Habr272727https://huggingface.co/datasets/IlyaGusev/habr and Habr QnA282828https://huggingface.co/datasets/its5Q/habr_qna are filtered by content popularity (e.g. views, ratings, votes). Cosine similarity obtained from LaBSE Feng et al. (2022) is applied to filter NewsCommentary and MultiParaCrawl. We filter pairs from paraphrase-NMT-Leipzig292929https://huggingface.co/datasets/cointegrated/ru-paraphrase-NMT-Leipzig by p_good score (equivalent meaning). The XNLI is formed from entailment (relevant document) and contradiction (irrelevant negative) examples. For MIRACL, we use the title as the query and the first paragraphs (until we reach the token limit) as the document. We form pairs for Paraphrases 303030https://huggingface.co/datasets/inkoziev/paraphrases from paraphrases field, taking one as a query and the others as positive documents. The content of RuNews313131https://huggingface.co/datasets/IlyaGusev/ru_news is not changed. After exact match deduplication, the final training pairs for all datasets are randomly sampled from the remaining pairs.
@@ A.2 Model Training Details
A.2.1 Default training details
We fine-tune the model in bf16 dtype with gradient checkpointing and use AdamW Loshchilov and Hutter (2017) with a learning rate of 1e-5 and weight decay of 0.01 for exactly one epoch, which is approximately 3700 steps, of which linear warmup is 200 steps. After fine-tuning, the SLERP merging is applied to the base model with a factor of 0.1.
We apply stratified sampling per device batch (mini-batch). Therefore, the global batch includes mini-batches consisting of pairs of different datasets. On the one hand, it becomes impossible to exchange negatives between devices and thus scale the number of in-batch negatives. On the other hand, this increases the diversity of sources in the global batch. Therefore, we do not apply the DisCo Chen et al. (2023) trick to exchange negative examples across devices. The batch size is 128 per device, giving 1024 documents per query. The context length is set to 512 for queries and documents Merrick et al. (2024).
Training is conducted on a single H100 node. We utilize the BGE323232https://github.com/FlagOpen/FlagEmbedding codebase and adapt it to our experiments. PyTorch’s expandable_segments helps us to mitigate fragmentation issues due to variable sequence length.
A.2.2 Ablation training details
Remove prefixes. We omit prefixes and keep the training process unchanged, preventing the model from identifying task types during training and inference.
Disable stratified sampling. Batch examples are randomly selected from all datasets instead of the single source. The prefixes are used only on the query side; otherwise the objective becomes easy to solve since different datasets have their own prefixes. DisCo is enabled to increase the number of negatives per query, unlike in stratified sampling.
Hard negatives. The number of hard negatives per query is increased from 7 to 15, while the batch size is reduced from 128 to 64 to maintain the same total number of negatives, and the gradient step accumulation (2 steps) is applied to keep training steps consistent.